2. Good Sleep Inc. is considering the purchase of a new mattress assembler that allows the construction of multiple firmness levels in the same mattress. Good Sleep anticipates the following 5-year earnings stream from the sale of this mattresS. Years in the Future Anticipated Profit (000s) 24 26 28 30 4 If the discount rate is 10% and the machine costs $90 (000), payable at once, what is the net present value of the machine? Is it a good purchase? (Write out the net earnings stream to generate your answer) Net Present Value Good purchase?:3. Joe Holiday has the opportunity to operate a business renting beach umbrellas next summer. He will operate the concession for 3 months. Looking at weather patterns, Joe observes that rain is frequent along this stretch of beach, and on average, there are 60 rentable days in a summer. In each of these days, Joe believes he can rent 40 umbrellas per day at $7 per rental. Joe will run the concession by himself day, and must pay Beachcomber Enterprises $9,000 for the concession (the use of the umbrellas and for the beachfront rental location) Suppose Joe could earn $4500 working construction. a. What are Joes Accounting Profits for undertaking the business? What are his Economic Profits Accounting profits (π Economic profits (): these profits attributable? Positive economic profits? (Y/N) Justification:
Answer #1

2.

R = 10%

Initial investment = $90 (,000)

Present value of benefits = 22/1.1 + 24/1.1^2 + 26/1.1^3 + 28/1.1^4 + 30/1.1^5

Present value of benefits = $97.12 (,000)

So,

Net present value = 97.12 - 90 = $7.12 (,000)

Good Purchase = Yes

It is a good purchase, because net present value is positive.

Pl. repost other unanswered questions for their proper answers!